# Learning to Evolve with Convergence Guarantee via Neural Unrolling

### v1

@@ VI-A Evolutionary Curriculum via Unshared Parameterization
VI-A1 Statistical Superiority of Unshared Design
The unshared parameterization strategy consistently outperforms the Shared baseline across all metrics. Table IV presents the aggregate performance across varying evaluation budgets. This advantage amplifies as the trajectory length increases. Specifically, at $\mathrm{MaxFEs}=20{,}000$ ($\mathrm{EMS}=200$), the Unshared model reduces the mean error by approximately 96% compared to the Shared version. This empirical gap suggests that a static operator is insufficient to handle the dynamic requirements of a complete optimization lifecycle. Conversely, the allocation of unique $\bm{\omega}$ to each block enables stage-wise specialization of the update dynamics. This capability is increasingly beneficial as $\mathrm{EMS}$ grows and the learning trajectory lengthens. Consequently, our final model adopts the separate per-block meta-parameters design for $\mathcal{O}_{\mathrm{evo}}$. Although this incurs a marginally higher training cost, it yields superior accuracy and stability under practical evaluation budgets.
@@ VI-A2 Stage-wise Specialization
Visual analysis confirms that the unshared parameterization enables an implicit evolutionary curriculum. Fig. 10 provides a visual explanation of this phenomenon. In the early stage, i.e., exploration, both configurations descend rapidly. However, in the late stage, i.e., exploitation, the Shared model (brown curves) flattens out prematurely. It functions as a stationary operator, which prevents mode switching. In contrast, the unshared model (red/purple curves) continues to refine the solution. The unshared parameterization effectively learns an implicit evolutionary curriculum. Specifically, early blocks specialize in global exploration to escape local basins, whereas later blocks shift towards precise local refinement. This time-dependent specialization is pivotal for high-precision convergence in black-box optimization.
@@ VI-B Deconstructing Component Synergy
We isolate the contributions of the three core architectural components of L2E: (i) ProxyGrad, which provides gradient injection; (ii) SoftGate, which enables adaptive fusion; and (iii) Mamba, which functions as the structured operator. We compare the full model, denoted as Ours, against variants where we individually remove or replace each component.
VI-B1 Quantitative Impact Analysis
Table V quantifies the performance degradation that results from the removal of individual components.
  • Impact of Gradient Guidance (w/o ProxyGrad): The removal of proxy gradients results in the most substantial performance deterioration, as the error increases by $\sim 103\times$. This outcome confirms that purely evolutionary search is sample-inefficient in high-dimensional spaces. Consequently, proxy gradients provide the descent momentum that is necessary to guide the population.
  • Impact of Adaptive Fusion (w/o SoftGate): The disablement of the gating mechanism, i.e., the enforcement of a deterministic combination, degrades performance by $\sim 161\times$. This result demonstrates the necessity of selective utilization. Specifically, the solver must dynamically decide when to trust the gradient for exploitation and when to rely on evolutionary proposals for exploration.
  • Impact of Architecture (w/o Mamba): The replacement of Mamba with a standard MLP/Attention block increases error by $\sim 6.8\times$. Although this degradation is less severe than the removal of gradients, it confirms that Mamba enables the modeling of long-range dependencies and coupled variables. This capacity is essential for breaking the precision ceiling on complex landscapes.
VI-B2 Convergence Dynamics
Fig. 11 and Fig. 12 visualize the per-problem behavior and convergence stability. On ill-conditioned functions, such as Ellipsoidal_hc, the Full model (red) achieves the lowest terminal cost and the tightest variance (Fig. 12, Right). In contrast, the w/o SoftGate variant exhibits high variance, which indicates that rigid updates lead to instability.
The proposed components function in concert. Specifically, ProxyGrad provides speed, Mamba ensures structural adaptation to landscape geometry, and SoftGate serves as a robust arbiter. The removal of any single component undermines this synergy, thereby leading to slower convergence or stagnation. Thus, L2E represents a cohesive system rather than a simple sum of parts. Its superior performance stems from the stage-wise curriculum, which is enabled by unshared unrolling, and the adaptive interplay between gradient-based and evolutionary search. This interplay is structurally supported by the non-expansive Mamba operator.
@@ VI-B2 Convergence Dynamics
Figure 12: Ablation on convergence and terminal cost across four variants (w/o ProxyGrad, w/o SoftGate, w/o Mamba, and Ours (Full)). Left: log objective gap vs. function evaluations (FEs) on two representative problems. Right: box-whisker statistics of terminal cost across runs.
@@ VII Conclusion
In this paper, we proposed L2E, a unified framework that approaches evolutionary optimization via neural unrolling. L2E enforces strict fixed-point contraction within a bilevel meta-learning scheme. This architecture seamlessly integrates the expressiveness of deep learning with the convergence guarantees of classical optimization. Empirical results across high-dimensional benchmarks and robot control tasks demonstrate the scalability and generalization capabilities. Limitations: However, the reliance on structured neural operators introduces computational overheads. Consequently, this computational cost restricts applicability in resource-constrained or low-latency environments. Future work will focus on developing lightweight neural operators that are compatible with real-time execution. Additionally, we aim to explore sample-efficient first-order algorithms to enhance inference speed. Finally, we plan to extend this non-expansive unrolling paradigm to multi-objective and constrained optimization scenarios.

### v2

@@ VI Ablation Analysis and Mechanistic Insights
VI-A1 Statistical Advantage of the Unshared Design
Across evaluation budgets, the unshared parameterization achieves consistently better performance than the shared-weight baseline. Table IV summarizes the aggregated results under different evaluation budgets, where the performance gap becomes more pronounced as the unrolling horizon increases. In particular, at $\mathrm{MaxFEs}=20{,}000$ ($\mathrm{EMS}=200$), the unshared model reduces the mean error by approximately 96% relative to the shared variant. These results suggest that a time-invariant update rule can be insufficient to accommodate the evolving needs of the optimization trajectory. By assigning step-specific meta-parameters $\bm{\omega}^{k}$, the unshared design enables stage-dependent update behaviors (e.g., exploration-oriented early steps and refinement-oriented later steps), which becomes increasingly beneficial as $\mathrm{EMS}$ grows. Therefore, we adopt separate per-block meta-parameters for $\mathcal{O}_{\mathrm{evo}}$ in the final model. Although this choice increases training cost moderately, it yields better accuracy and more stable behavior under practical evaluation budgets.
@@ VI-A2 Stage-wise Specialization
Figure 10 provides qualitative evidence that unshared parameterization enables stage-dependent update behaviors along the unrolled trajectory. In the early phase (exploration), both the shared and unshared designs make rapid progress. However, in the late phase (refinement), the shared-weight variant (brown curves) tends to plateau earlier, indicating limited flexibility to adapt its update behavior as the search proceeds. In contrast, the unshared variant (red/purple curves) continues to refine the solutions and achieves better late-stage improvements. Overall, this visualization suggests that step-specific meta-parameters $\bm{\omega}^{k}$ allow different unrolling blocks to specialize in different roles with earlier blocks favoring broader exploration and later blocks emphasizing local refinement, which is beneficial for achieving higher final precision in black-box optimization.
@@ VI-B Deconstructing Component Synergy
We further isolate the contributions of three core components in L2E: (i) ProxyGrad, which injects proxy numerical guidance; (ii) SoftGate, which enables adaptive fusion; and (iii) Mamba, which parameterizes the structured evolutionary operator. We compare the full model (Ours) with variants that remove or replace each component to quantify their individual and combined effects.
@@
VI-B1 Quantitative Impact Analysis
Table V quantifies the performance changes when individual components are removed or replaced.
  • Impact of Proxy Guidance (w/o ProxyGrad): Removing proxy gradient guidance leads to the largest degradation, increasing the error by $\sim 103\times$. This result indicates that, under the considered evaluation budgets, purely evolutionary updates can be substantially less sample-efficient, whereas proxy guidance provides an effective directional signal for accelerating progress.
  • Impact of Adaptive Fusion (w/o SoftGate): Disabling the gating mechanism (i.e., using a fixed combination) degrades performance by $\sim 161\times$. This result underscores the necessity of selective fusion. Specifically, the solver benefits from dynamically balancing proxy guidance for refinement and learned evolutionary proposals for exploration. In contrast, a rigid mixing rule lacks this essential flexibility.
    Figure 12: Ablation on convergence and terminal cost across four variants. Left: log objective gap vs. function evaluations (FEs) on two representative problems. Right: box-whisker statistics of terminal cost across runs.
  • Impact of Operator Parameterization (w/o Mamba): Replacing the Mamba-based operator with a standard MLP/attention block increases the error by $\sim 6.8\times$. Although this degradation is smaller than that caused by removing guidance or gating, it indicates that the structured operator parameterization enhances the modeling of population-wise interactions and variable couplings. This capability is particularly beneficial on complex landscapes.
VI-B2 Convergence Dynamics
Figure 11 and Fig. 12 further visualize convergence behavior and run-to-run variability. On ill-conditioned functions such as Ellipsoidal_hc, the full model (red) achieves better terminal performance with tighter variance (Fig. 12, Right). In contrast, the w/o SoftGate variant exhibits noticeably higher variance, consistent with the intuition that rigid update rules can be brittle across diverse instances. Overall, the three components act in a complementary manner. ProxyGrad provides effective numerical guidance, Mamba offers a structured parameterization for population updates, and SoftGate adaptively mediates between the two update sources. Removing any single component weakens this complementarity and often leads to slower progress or earlier stagnation. Together with unshared unrolling parameters, these design choices promote stage-dependent behaviors along the trajectory and enable stable, effective optimization across diverse numerical landscapes.
@@ VII Conclusion
In this paper, we proposed L2E, a unified framework that approaches evolutionary optimization via neural unrolling. Instead of learning unconstrained update rules, L2E adopts a stability-inducing unrolling formulation within a bilevel meta-learning scheme, which promotes controlled optimization dynamics in practice. Empirical results on high-dimensional numerical benchmarks and robot control tasks demonstrate the scalability of L2E and its robust performance under distribution shifts. Limitations: The framework introduces additional training-time overhead due to unrolling and structured operator parameterization. This cost may limit applicability in resource-constrained or low-latency settings. Future Work: We will develop lightweight operator parameterizations and more efficient training strategies to reduce computation and memory usage. We also plan to extend L2E to broader settings, such as constrained and multi-objective optimization.